In September direct labor was 35% of conversion cost. If the manufacturing overhead for the month was $102,050 and the direct materials cost was $23,800, the direct labor cost was:

A manufacturing company prepays its insurance coverage for a three-year period. The premium for the three years is $3,600 and is paid at the beginning of the first year. Eighty percent of the premium applies to manufacturing operations and twenty percent applies to selling and administrative activities. What amounts should be considered product and period costs respectively for the first year of coverage?
Answer #1

Conversion cost = Direct labor cost + manufacturing overhead cost

Let the conversion cost be x

X = (35%*x + 102,050)

X = $157,000

Direct labor cost = $54,950

Manufacturing costs are a part of product cost

Product cost during first year = 3600*1/3*80% = $960

Selling and admin costs are period costs i.e. 1200*20% = $240